OpenStudy (anonymous):

Consider the line that passes through the points (4, -3) and (2, -1). Below are two different equations. Identify the true statement concerning both equations. Equation #1 Equation #2 y - 3 = -1(x + 4) y - 1 = -1(x + 2) Only equation #1 represents the line that passes through the two given points. Only equation #2 represents the line that passes through the two given points. Both equation #1 and equation #2 represent the line that passes through the two given points. Neither equation #1 nor equation #2 represents the line that passes through the two

OpenStudy (anonymous):

try finding the slope of the line that passes between the two points,(4, -3) and (2, -1). Once you have found the slope, check equation 1 and 2 to see if either has that same slope.

OpenStudy (anonymous):

how would i find the slope

OpenStudy (anonymous):

slope is (-3 -(-1) / (4 - 2) change in the y values divided by change in the x values. Another different approach is to take the 2 points one at a time, and to try the x and y values from each point as inputs into each line. If you put the values in and the equation works, it means that point is on the line. If you put the points in and you get nonsense, then the point is not on the line.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Equation #1 Equation #2 y - 3 = -1(x + 4) y - 1 = -1(x + 2) 1 - 3 = -1(-3 + 4) 2 - 1 = -1(7 + 2)

OpenStudy (anonymous):

Are you seeing "nonsense" in those equations?

OpenStudy (anonymous):

#1 1 - 3 = -1(-3 + 4) -2 = -3 + -4

OpenStudy (anonymous):

-2 = -3 - 4 +4 +4 2 = -3 i see nonsense in #1

OpenStudy (anonymous):

hold on... you mixed it up a little, although you are good at spotting nonsense :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Let's try Eq # 1 again, using the 2nd point, (2, -1): y - 3 = -1(x + 4) (-1) - 3 = -1(2 + 4) -4 = -1(6) -4 = -6 nonsense... therefore, point (2,-1) is NOT on the line shown in Eq # 1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

if you wanted, you could also try the first point, (4, -3) on the same Eq # 1, but it doesn't matter for Eq # 1... if the second point isn't even on the line, it doesn't matter if the other point is on the line... the line itself doesn't go through both points. We still need to check Eq # 2

OpenStudy (anonymous):

Equation #2 y - 1 = -1(x + 2) 2 - 1 = -1(-1 + 2) 3 = 1 - 2 +2 +2 5 = 1 Nonsense

OpenStudy (anonymous):

again, close, but be careful... you chose point (2, -1), but when you put it in the equation, you substituted in the (-1) term as the "x" on the right side... the point is (x,y), so you need to put in x = 2 on the right side and y = -1 on the left. It will probably still be nonsense, but please be careful substituting in the (x,y)

OpenStudy (anonymous):

ok but is the answer D???